OpenStudy (anonymous):

will medal and fan multiply and simplify 6 1/4 and 4/5

OpenStudy (jdoe0001):

\(\bf 6\cfrac{1}{4}\cdot \cfrac{4}{5}\implies \cfrac{6\cdot {\color{blue}{ 4}}+1}{{\color{blue}{ 4}}}\cdot \cfrac{4}{5}\implies \cfrac{(6\cdot {\color{blue}{ 4}}+1)\cdot 4}{4\cdot 5}\)

OpenStudy (anonymous):

@sammixboo could come to my queastion please

OpenStudy (anonymous):

so wats the answer @jdoe0001

sammixboo (sammixboo):

Okie Dokie. We can solve this by following the order of operations. P - Parenthesis E - Exponents M/D - Multiplication or division A/S - Addition or subtraction As we can see, there is parenthesis. Let's solve what is in the parenthesis first. \(\Large\frac{(6\times4~+~1)4}{4\times5}\) We can see that 6 x 4 + 1 is in the parenthesis. We must multiply first, so what is 6 x 4?

OpenStudy (anonymous):

24

sammixboo (sammixboo):

Good! Now we have \(\Large\frac{(24~+~1)4}{4\times5}\) Still solving what is in the parenthesis, all we have is 24 + 1, which is 25, so we plug in 25 where (24+1) was at So now we have \(\Large\frac{(25)4}{4\times5}\) Can you tell me what is 25x4 is?

OpenStudy (anonymous):

so the mph is like what

OpenStudy (anonymous):

100

OpenStudy (anonymous):

@sammixboo